# FakET: Simulating Cryo-Electron Tomograms with Neural Style Transfer

### v3

@@ Abstract
Correspondence to— P. Harar <<pavol.harar@univie.ac.at>> and D. Haselbach <<david.haselbach@imp.ac.at>>
@@ 1 Introduction
In cryoET, the imaged sample is in most cases a 100-200 nm thick slice of a frozen cell. From this slice, projection images are taken in a transmission electron microscope (TEM) from different rotation (tilt) angles. An artifact free reconstruction would require measurements using tilt angles, that would complete the half circle. However, this is not feasible, due to limitations of the specimen holder, and only a range of 140 can be recorded. The missing tilt images later on result in a so-called missing-wedge in the 3D reconstruction (cryo-electron tomogram). In addition, the electron beam severely damages the sample during imaging, so only a low electron dose can be used to image a biological specimen. The low dose in combination with the presence of ice in the sample results in the acquired data being very noisy. Consequently, the identification of molecules within these reconstructions is a daunting task. Particle identification is however necessary as the particles need to be classified and averaged to determine high resolution structures. While cryoET has led to a large number of breakthroughs, providing hitherto unseen detail in the molecular architecture of cells (Zimmerli et al., 2021; O’Reilly et al., 2020; Mahamid et al., 2016), the aforementioned challenges still hinder the widespread use of cryoET in the larger cell biology and structural biology community. In this context, the development of new reliable software tools is of paramount importance, which is however obstructed by the lack of sufficient accessible and annotated data to develop the software tools on.
@@ 1.1 SHREC simulator
To overcome the problem with the lack of data, in 2019, the annual SHREC – 3D Shape Retrieval Contest included a new track titled Classification in Cryo-Electron Tomograms. The organizers of this track proposed a task of localization and classification of biological particles in cryo-electron tomograms. In the following years, experts from 3D object retrieval and 3D electron microscopy communities were invited to participate in the challenge. In order to ensure fair evaluation and comparable results across the submissions, the organizers created a data set of ten physics-based cryo-electron tomogram simulations (9 train & 1 test tomogram) for the contestants to train and evaluate their methods on. Each year, the results of the contesting methods were presented and compared Gubins et al. (2019, 2020, 2021).
@@ 1.2 Our contribution
In this paper, we propose faket, a fast and scalable data-driven method for simulating the forward operator of any cryo transmission electron microscope with the aim to generate synthetic micrographs or tilt-series. It was created, among other reasons, to generate fully-labeled real-like data for training deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. Our method combines additive noise and neural style transfer (NST) technique based on Gatys et al. (2016) to capture, using only unlabeled reference data, the structure of the complex noise introduced by TEM. To carry out the NST, faket utilizes a pretrained model, eliminating the need for users to train it themselves. That means users can use faket to simulate data from any TEM and under any configuration, assuming they possess unlabeled TEM data that could serve as a style reference. faket delivers data of quality nearly identical [^1] In terms of their practical utility for subsequent tasks, rather than exact manifestation of the physical interactions between the sample and the transmission electron microscope. to the reference, at a cost only slightly higher than simply adding Gaussian noise to a noiseless synthetic sample [^2] Created using existing models of biological macromolecular structures, represented as Coulomb density volume.. This makes it a practical, effective, and efficient choice for simulation.
@@
In contrast to SHREC, our method accelerates the data generation process by a factor of 750 while using 33 times less memory (see Table 1). It also does not require any calibration protocol as other simulators, see Section 1.4. The NST model does not need to be retrained to be used on new data, nor does it require labeled reference data. faket therefore has the potential to save experts countless hours [^4] Depending on the imaged sample, an expert may spend several hours of manual work per tilt-series per particle labeling the data. At the same time, certain smaller particles cannot be visually found at all, thus fully labeled tomograms do not really exist. of manual work in labeling their data sets. Moreover, faket is capable of simulating large tilt-series, which are common in experimental environments (about $12\times$ larger than SHREC projections). For example, we generated a $61\times 3500\times 3500$ tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes [^5] To achieve the reported speeds, a minimum of one GPU is recommended; CPU-only simulations are feasible and can be a valid option in specific situations, albeit slower. This advance makes it possible to train particle localization and classification networks from scratch, or to pre-train networks that are later fine-tuned using manually labelled experimental data, see Appendix C. On top of that, our method is open source and our experiments are reproducible.
Additionally, we include comparisons of the benchmark and the proposed faket method with a simple baseline method based on additive Gaussian noise aiming at providing comparative results in full practical range. The costly SHREC method defines an upper-boundary of performance, while the cheap addition of Gaussian noise establishes a lower boundary. This configuration enables us to position our method along this spectrum, showcasing faket’s capability to offer results close to the upper-boundary while maintaining computational efficiency of a much simpler method. We hope our comparisons will be useful for practitioners who need to decide which method fits into their computational budget and for those seeking insightful understanding of the inherent trade-offs. The contribution of our method is further supported by an ablation study presented in Appendix D, where we also offer insight into the potential limits of DF on the studied data by conducting experiments using completely noiseless simulated tomograms.
@@ 1.4 Other simulators
Earlier works provided fast simplistic models mostly based on additive white Gaussian noise [^7] This motivated our choice of baseline. or coloured noise Scheible et al. (2021). Other works originated from the insights into the physics of TEM image formation and advanced the simulations by modeling various sources of noise, e.g. an improvement in modeling the structural noise was proposed in TEM Simulator (C) Rullgård et al. (2011). Another improvement was done in InSilicoTEM (Matlab) presented in Vulović et al. (2013) by implementing the multislice method originally proposed in Cowley and Moodie (1957). This makes it the next most relevant simulator related to our work after SHREC which is, as many others, also based on the same multislice method  [^8] This motivated our choice of benchmark.. To the best of our knowledge, the most recent improvement presented in Himes and Grigorieff (2021) is distributed as a part of cisTEM package (C++). It introduced frozen plasmon method to explicitly model spatially variable inelastic scattering processes in cryo-electron microscopy. The aforementioned works, based on the same multislice method, however, suffer from a heavy computational burden or are limited to simulations of a single molecular complex. Despite these limitations, they represent a set of precise advanced physics-based simulators and as such also provide a comprehensive literature survey referring the reader to a body of detailed resources about modeling the image formation in transmission electron microscopy.
@@ 1.5 Structure of this article
To evaluate the methods, we used the simulated data by each of them as a training set for the DeepFinder neural network (more in Section 4.1). The network was trained to solve two evaluation tasks proposed in SHREC challenge. The final comparison was done by observing the models’ performances on the test tomogram chosen by SHREC. Detailed description of the evaluation is in Section 4. More on experiments and results in Section 5. This paper is accompanied with source code, the results are fully-reproducible, and the full experiment results are available in the repository.
@@ 2 Data
For the purposes of this paper, we used the latest version of the SHREC data set from the year 2021. More specifically, we based the presented methods on full-resolution grandmodels (synthetic volumes containing randomly scattered particles) and simulated projections from the SHREC 2021 additional data set that was made publicly available [^9] Download from shrec.net/cryo-et or doi:10.34894/XRTJMA.. An example visualization of a simulated projection using the SHREC simulator is depicted in Figure 1. We use the unchanged simulated projections as a benchmark which we try to approximate with our proposed method.
@@
Figure 1: Simulated projection from SHREC 2021 data set. Axes x and y correspond to width and height of the imaged grandmodel. Colorbar denotes simulated intensities in arbitrary units. See Appendix A for side-by-side comparison with other projections.
@@ 3 Methods
Figure 3: baseline projections. Axes x and y correspond to width and height of the imaged grandmodel. Colorbar denotes simulated intensities in arbitrary units. Please note that this projection is not exactly the same as in Figure 1 or in Figure 4, see also the explanation in the caption of Figure 4 and Appendix A for a side-by-side comparison with other projections.
@@ 3.1 Additive noise (baseline)
Figure 4: faket projections. Axes x and y correspond to the grandmodel’s width and height respectively, and colorbar indicates simulated intensities in arbitrary units. While this projection may appear visually similar to Figure 3, there is a subtle difference that has a significant impact on DF’s performance. This similarity complicates comparison of projections from various simulators using currently available image metrics. See Appendix A for a side-by-side comparison with other projections.
@@ 5 Results
Figure 6: Performance on localization task.
@@
To put these numbers in perspective, we investigated the limits of DF performance by training and evaluating on completely noiseless data, see Appendix D for more details. In this setting DF achieved F1 score of 83.2 % (68 % CI [82.4 %, 84.0 %]) on localization task and 72.5 % (68 % CI [70.5 %, 74.6 %]) on classification task.
@@
Figure 7: Performance on classification task.
Figure 8: Per-class performance on classification task.
@@ Appendix A Side-by-side comparison of projection types
Figure 9: Side-by-side comparison of all different types of projections presented in this article.
Figure 10: One instance of style projections (from benchmark) used as a reference for NST (left) and the difference between one particular instance of faket and benchmark projections. Note that the style projection used as a reference input to NST never features the same content as the input projection that is being style-transferred. Additional comparisons of projections through visualizations and standard metrics are presented in the supplement.
@@ Appendix B Recreating benchmark – reverse-engineering SHREC filtering step
In pursuit of re-creating the benchmark reconstructions to match those from SHREC data set as closely as possible, we inspected the public source code of PyTom Hrabe et al. (2012) for clues on which filters might have been used. In the original paper, authors only mention that they use weighted back-projection. Which would suggest the use of a simple ramp filter. In a personal communication, they mentioned using a default filter from their private PyTom version with undefined Crowther frequency. Therefore, we implemented and used a simple ramp filter with a support for Crowther frequency. Then, we visualized SHREC’s and our reconstructions in Fourier space to visually compare the results. Unfortunately, none of the variations of a simple ramp filter, even with Crowther frequency, produced the expected result similar to SHREC. For that reason, we started to iteratively change our filter in order to approximately match the aforementioned reconstructions in both real and Fourier space, in hope of finding the desired match. After many iterations, we found a good-enough match, a filter which is a product of a 2D Gaussian filter ($\sigma_{x}=174$, $\sigma_{y}=102$), a 1D ramp filter broadcasted to 2D ($\mathrm{CrowtherFreq}=0.61$ Nyquist), and a 2D circular filter ($\mathrm{radiusCutoff}=256$, i.e. Nyquist frequency). Exact implementation is available in code repository.
@@
Figure 11: Middle Z-axis slice of SHREC (left) and benchmark (right) reconstructions.
@@
Here, we compared the performance of models trained on baseline or faket data with that of a model initially trained on faket data and later fine-tuned using a subset (3 out of 9) of benchmark tomograms, denoted as finetuned-3 in the figures. The fine-tuning began after the 50th epoch, triggering a notable improvement in both localization and classification performance. To make a reasonable comparison, we include results from a model, benchmark-3, trained solely on the same 3 benchmark tomograms used for fine-tuning.
@@
Figure 12: Performance of finetuned-3 and benchmark-3 models in context with the main results.
@@ Appendix D Additional experiments
Figure 13: Performance of DF trained with noisy (top), faket-random (middle), and noiseless (bottom) data. Localization results are shown in the left column and classification results in the right column.
@@ Appendix E Further details on particle classification performance
Next to F1 scores reported in Section 5, here we present aggregated confusion matrices for all main and selected additional experiments. Each confusion matrix represents the average of 6 confusion matrices obtained from 6 different experiments each initialized with a different random seed. Each confusion matrix was computed based on the results of the best epoch that was selected based on the model’s validation performance. The csv files containing information about which epochs were selected for each of the aggregated confusion matrices are available in the accompanying repository.
@@
Figure 14: Aggregated confusion matrices for benchmark (top) and faket (bottom) experiments.
Figure 15: Aggregated confusion matrices for baseline (top) and noiseless (bottom) experiments.
Figure 16: Aggregated confusion matrices for finetuned-3 (top) and benchmark-3 (bottom) experiments.
@@ Appendix F Profiling execution time and memory consumption
Profiling was done using an NVIDIA A100 40GB SXM4 GPU, a Dual AMD EPYC 7742 64-Core CPU with total 128 physical CPU cores and 256 threads, a DDR4-3200 MHz memory, and Samsung PM983 NVMe read-intensive hard drives.

### v4

@@ Abstract
Correspondence to— P. Harar <<pavol.harar@ista.ac.at>> and D. Haselbach <<david.haselbach@imp.ac.at>>
@@ 1 Introduction
In cryoET, the imaged sample is in most cases a 100-200 nm thick slice of a frozen cell. From this slice, projection images are taken in a transmission electron microscope (TEM) from different rotation (tilt) angles. An artifact free reconstruction would require measurements using tilt angles, that would complete the half circle. However, this is not feasible, due to limitations of the specimen holder, and only a range of 140 can be recorded. The missing tilt images later on result in a so-called missing-wedge in the 3D reconstruction (cryo-electron tomogram). In addition, the electron beam severely damages the sample during imaging, so only a low electron dose can be used to image a biological specimen. The low dose in combination with the presence of ice in the sample results in the acquired data being very noisy. Consequently, the identification of molecules within these reconstructions is a daunting task. Particle identification is however necessary as the particles need to be classified and averaged to determine high resolution structures. While cryoET has led to a large number of breakthroughs, providing hitherto unseen detail in the molecular architecture of cellsZimmerli et al. (2021); O’Reilly et al. (2020); Mahamid et al. (2016), the aforementioned challenges still hinder the widespread use of cryoET in the larger cell biology and structural biology community. In this context, the development of new reliable software tools is of paramount importance, which is however obstructed by the lack of sufficient accessible and annotated data to develop the software tools on.
@@ 1.1 SHREC simulator
To overcome the problem with the lack of data, in 2019, the annual SHREC – 3D Shape Retrieval Contest included a new track titled Classification in Cryo-Electron Tomograms. The organizers of this track proposed a task of localization and classification of biological particles in cryo-electron tomograms. In the following years, experts from 3D object retrieval and 3D electron microscopy communities were invited to participate in the challenge. In order to ensure fair evaluation and comparable results across the submissions, the organizers created a data set of ten physics-based cryo-electron tomogram simulations (9 train & 1 test tomogram, see Figure 1) for the contestants to train and evaluate their methods on. Each year, the results of the contesting methods were presented and comparedGubins et al. (2019, 2020, 2021).
@@ 1.2 Our contribution
In this paper, we propose faket, a fast and scalable data-driven method for simulating the forward operator of any cryo transmission electron microscope with the aim to generate synthetic micrographs or tilt-series. It was created, among other reasons, to generate fully-labeled real-like data for training deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. Our method combines a noiseless simulated sample, see Figure 2, with additive noise, see Figure 3, and neural style transfer (NST) technique based onGatys et al. (2016) to capture, using only unlabeled reference data, the structure of the complex noise introduced by TEM, see Figure 4. To carry out the NST, faket utilizes a pretrained model, eliminating the need for users to train it themselves. That means users can use faket to simulate data from any TEM and under any configuration, assuming they posses unlabeled TEM data that could serve as a style reference. faket delivers data of quality nearly identical [^1] In terms of their practical utility for subsequent tasks, rather than exact manifestation of the physical interactions between the sample and the transmission electron microscope. to the reference, at a cost only slightly higher than simply adding Gaussian noise to a noiseless synthetic sample [^2] Created using existing models of biological macromolecular structures, represented as Coulomb density volume.. This makes it a practical, effective, and efficient choice for simulation.
@@
In contrast to SHREC, our method accelerates the data generation process by a factor of 750 while using 33 times less memory (see Table 1). It also does not require any calibration protocol as other simulators, see Section 1.4. The NST model does not need to be retrained to be used on new data, nor does it require labeled reference data. faket therefore has the potential to save experts countless hours of manual work in labeling their data sets. [^4] Depending on the imaged sample, an expert may spend several hours of manual work per tilt-series per particle labeling the data. At the same time, certain smaller particles cannot be visually found at all, thus fully labeled tomograms do not really exist. Moreover, faket is capable of simulating large tilt-series, which are common in experimental environments (about $12\times$ larger than SHREC projections). For example, we generated a $61\times 3500\times 3500$ tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. [^5] To achieve the reported speeds, a minimum of one GPU is recommended; CPU-only simulations are feasible and can be a valid option in specific situations, albeit slower. This advance makes it possible to train particle localization and classification networks from scratch, or to pre-train networks that are later fine-tuned using manually labeled experimental data, see Appendix C. On top of that, our method is open source and our experiments are reproducible.
Additionally, we include comparisons of the benchmark and the proposed faket method with a simple baseline method based on additive Gaussian noise, see Figure 3, aiming at providing comparative results in full practical range. The costly SHREC method defines an upper-boundary of performance, while the cheap addition of Gaussian noise establishes a lower boundary. This configuration enables us to position our method along this spectrum, showcasing faket’s capability to offer results close to the upper-boundary while maintaining computational efficiency of a much simpler method. We hope our comparisons will be useful for practitioners who need to decide which method fits into their computational budget and for those seeking insightful understanding of the inherent trade-offs. The contribution of our method is further supported by an ablation study presented in Appendix D, where we also offer insight into the potential limits of DF on the studied data by conducting experiments using completely noiseless simulated tomograms.
@@ 1.4 Other simulators
Earlier works provided fast simplistic models mostly based on additive white Gaussian noise [^7] This motivated our choice of baseline. or coloured noiseScheible et al. (2021). Other works originated from the insights into the physics of TEM image formation and advanced the simulations by modeling various sources of noise, e.g. an improvement in modeling the structural noise was proposed in TEM Simulator (C)Rullgård et al. (2011). Another improvement was done in InSilicoTEM (Matlab) presented inVulović et al. (2013) by implementing the multislice method originally proposed inCowley and Moodie (1957). This makes it the next most relevant simulator related to our work after SHREC which is, as many others, also based on the same multislice method. [^8] This motivated our choice of benchmark. To the best of our knowledge, the most recent improvement presented inHimes and Grigorieff (2021) is distributed as a part of cisTEM package (C++). It introduced frozen plasmon method to explicitly model spatially variable inelastic scattering processes in cryo-electron microscopy. The aforementioned works, based on the same multislice method, however, suffer from a heavy computational burden or are limited to simulations of a single molecular complex. Despite these limitations, they represent a set of precise advanced physics-based simulators and as such also provide a comprehensive literature survey referring the reader to a body of detailed resources about modeling the image formation in transmission electron microscopy.
@@ 1.5 Structure of this article
To evaluate the methods, we used the simulated data by each of them as a training set for the DeepFinder neural network (more in Section 4.1). The network was trained to solve two evaluation tasks proposed in SHREC challenge. The final comparison was done by observing the models’ performances on the test tomogram chosen by SHREC. Detailed description of the evaluation is in Section 4. More on experiments and results in Section 5. Procedures of each method are compared in Figure 5. This paper is accompanied with source code, the results are fully-reproducible, and the full experiment results are available in the repository – https://github.com/paloha/faket/.
@@ 2 Data
For the purposes of this paper, we used the latest version of the SHREC data set from the year 2021. More specifically, we based the presented methods on full-resolution grandmodels (synthetic volumes containing randomly scattered particles) and simulated projections from the SHREC 2021 additional data set that was made publicly available [^9] Download v2.0: shrec.net/cryo-et, doi:10.34894/XRTJMA.. An example visualization of a simulated projection using the SHREC simulator is depicted in Figure 1. We use the unchanged simulated projections as a benchmark which we try to approximate with our proposed method.
@@
Figure 1: Simulated projection taken from SHREC 2021 data set. Axes x and y correspond to width and height of the imaged grandmodel. Colorbar denotes simulated intensities in arbitrary units. See Appendix A for side-by-side comparison with other projections.
@@ 3 Methods
Figure 3: baseline projection created by adding Gaussian noise to the noiseless projection. Axes x and y correspond to width and height of the imaged grandmodel. Colorbar denotes simulated intensities in arbitrary units. Please note that this projection is not exactly the same as in Figure 1 or in Figure 4, see also the explanation in the caption of Figure 4 and Appendix A for a side-by-side comparison with other projections.
@@ 3.1 Additive noise (baseline)
Figure 4: faket projection output by our method. Axes x and y correspond to the grandmodel’s width and height respectively, and colorbar indicates simulated intensities in arbitrary units. Though visually similar to Figure 3, there is a subtle difference that has a significant impact on DF’s performance. This similarity complicates comparison of projections from various simulators using currently available image metrics. See Appendix A for a side-by-side comparison with other projections.
@@ 5 Results
Figure 6: Performance on localization task. Darker and lighter regions show 68% and 95% CI.
@@
Figure 7: Performance on classification task. Darker and lighter regions show 68% and 95% CI.
@@
Figure 8: Per-class performance on classification task. Darker and lighter regions show 68% and 95% CI.
@@ Appendix A Side-by-side comparison of projection types
Figure 9: Side-by-side comparison of all types of projections presented in this article.
Figure 10: One instance of style projections (from benchmark) used as a reference for NST (left) and the difference between one particular instance of faket and benchmark projections. Note that the style projection used as a reference input to NST never features the same content as the input projection that is being style-transferred. Additional comparisons of projections through visualizations and standard metrics are presented in the supplement. Colorbar indicates simulated intensities in arbitrary units.
@@ Appendix B Recreating benchmark – reverse-engineering SHREC filtering step
In pursuit of re-creating the benchmark reconstructions to match those from SHREC data set as closely as possible, we inspected the public source code of PyTomHrabe et al. (2012) for clues on which filters might have been used. In the original paper, authors only mention that they use weighted back-projection. Which would suggest the use of a simple ramp filter. In a personal communication, they mentioned using a default filter from their private PyTom version with undefined Crowther frequency. Therefore, we implemented and used a simple ramp filter with a support for Crowther frequency. Then, we visualized SHREC’s and our reconstructions in Fourier space to visually compare the results. Unfortunately, none of the variations of a simple ramp filter, even with Crowther frequency, produced the expected result similar to SHREC. For that reason, we started to iteratively change our filter in order to approximately match the aforementioned reconstructions in both real and Fourier space, in hope of finding the desired match, see Figure 11. After many iterations, we found a good-enough match, a filter which is a product of a 2D Gaussian filter ($\sigma_{x}=174$, $\sigma_{y}=102$), a 1D ramp filter broadcasted to 2D ($\mathrm{CrowtherFreq}=0.61$ Nyquist), and a 2D circular filter ($\mathrm{radiusCutoff}=256$, i.e. Nyquist frequency). Exact implementation is available in code repository.
@@
Figure 11: Middle Z-axis slice of SHREC (left) and benchmark (right) reconstructions. Colorbar indicates simulated intensities in arbitrary units (top), and absolute value of the Fourier transform (bottom).
@@
Here, we compared the performance of models trained on baseline or faket data with that of a model initially trained on faket data and later fine-tuned using a subset (3 out of 9) of benchmark tomograms, denoted as finetuned-3 in the Figure 12. The fine-tuning began after the 50th epoch, triggering a notable improvement in both localization and classification performance. To make a reasonable comparison, we include results from a model, benchmark-3, trained solely on the same 3 benchmark tomograms used for fine-tuning.
@@
Figure 12: Performance of finetuned-3 and benchmark-3 models in context with the main results. Darker and lighter regions show 68% and 95% CI.
@@ Appendix D Additional experiments
Figure 13: Performance of DF trained with noisy (top), faket-random (middle), and noiseless (bottom) data. Results of localization (left) and classification (right). Darker and lighter regions show 68% and 95% CI.
@@ Appendix E Further details on particle classification performance
Next to F1 scores reported in Section 5, here we present aggregated confusion matrices for all main particle classification and selected additional experiments, see Figures 14, 15 and 16. Each confusion matrix represents the average of 6 confusion matrices obtained from 6 different experiments each initialized with a different random seed. Each confusion matrix was computed based on the results of the best epoch that was selected based on the model’s validation performance. The csv files containing information about which epochs were selected for each of the aggregated confusion matrices are available in the accompanying repository.
@@
Figure 14: Confusion matrices for benchmark (top) and faket (bottom) experiments.
Figure 15: Confusion matrices for baseline (top) and noiseless (bottom) experiments.
Figure 16: Confusion matrices for finetuned-3 (top) and benchmark-3 (bottom) experiments.
@@ Appendix F Profiling execution time and memory consumption
Profiling of FakET was done using an NVIDIA A100 40GB SXM4 GPU, a Dual AMD EPYC 7742 64-Core CPU with total 128 physical CPU cores and 256 threads, a DDR4-3200 MHz memory, and Samsung PM983 NVMe read-intensive hard drives.
@@
Supplementary Information
model modality
MAE mean
MAE $10^{th}$
MAE $90^{th}$
MSE mean
MSE $10^{th}$
MSE $90^{th}$
PSNR mean
PSNR $10^{th}$
PSNR $90^{th}$
SSIM mean
SSIM $10^{th}$
SSIM $90^{th}$
all faket 0.37 0.32 0.45 0.15 0.11 0.21 8.40 6.80 9.57 0.25 0.17 0.38
noisy 0.38 0.35 0.42 0.15 0.13 0.18 8.16 7.44 8.86 0.19 0.17 0.21
shrec 0.53 0.47 0.61 0.29 0.23 0.38 5.45 4.22 6.44 0.17 0.13 0.21
0 faket 0.38 0.34 0.42 0.16 0.12 0.18 8.14 7.38 9.12 0.19 0.17 0.20
noisy 0.41 0.39 0.43 0.18 0.16 0.19 7.47 7.20 7.97 0.18 0.16 0.19
shrec 0.60 0.53 0.64 0.37 0.29 0.42 4.34 3.79 5.37 0.13 0.13 0.14
1 faket 0.36 0.35 0.39 0.14 0.13 0.16 8.56 8.08 8.87 0.18 0.16 0.22
noisy 0.37 0.34 0.38 0.14 0.12 0.15 8.54 8.25 9.07 0.20 0.18 0.21
shrec 0.51 0.47 0.55 0.27 0.23 0.31 5.67 5.12 6.41 0.17 0.15 0.17
2 faket 0.33 0.30 0.35 0.11 0.10 0.12 9.47 9.07 10.20 0.38 0.36 0.41
noisy 0.38 0.36 0.39 0.15 0.13 0.16 8.22 7.94 8.75 0.19 0.18 0.20
shrec 0.53 0.52 0.55 0.29 0.28 0.31 5.37 5.02 5.58 0.13 0.11 0.16
3 faket 0.34 0.31 0.36 0.12 0.10 0.13 9.19 8.79 9.89 0.36 0.33 0.38
noisy 0.39 0.37 0.41 0.16 0.14 0.17 7.87 7.59 8.41 0.17 0.15 0.17
shrec 0.48 0.44 0.50 0.24 0.20 0.26 6.28 5.88 6.98 0.20 0.19 0.21
4 faket 0.38 0.33 0.45 0.15 0.12 0.21 8.29 6.88 9.22 0.20 0.17 0.22
noisy 0.37 0.35 0.39 0.14 0.13 0.15 8.40 8.13 8.92 0.20 0.18 0.21
shrec 0.48 0.44 0.50 0.24 0.20 0.26 6.27 5.87 6.97 0.21 0.20 0.22
5 faket 0.36 0.33 0.38 0.13 0.11 0.15 8.73 8.34 9.45 0.34 0.33 0.37
noisy 0.41 0.38 0.42 0.17 0.15 0.18 7.62 7.36 8.13 0.18 0.17 0.19
shrec 0.52 0.48 0.60 0.29 0.24 0.37 5.49 4.36 6.23 0.14 0.13 0.15
6 faket 0.38 0.32 0.41 0.15 0.11 0.17 8.31 7.59 9.56 0.24 0.22 0.26
noisy 0.35 0.33 0.36 0.13 0.11 0.14 8.93 8.65 9.47 0.21 0.19 0.22
shrec 0.50 0.46 0.52 0.26 0.22 0.28 5.93 5.52 6.65 0.19 0.18 0.20
7 faket 0.43 0.39 0.50 0.19 0.16 0.25 7.23 5.99 8.05 0.19 0.17 0.22
noisy 0.39 0.37 0.41 0.16 0.14 0.17 7.93 7.66 8.45 0.19 0.17 0.20
shrec 0.57 0.51 0.60 0.33 0.26 0.37 4.89 4.37 5.78 0.17 0.17 0.18
8 faket 0.42 0.34 0.47 0.19 0.13 0.22 7.36 6.49 8.99 0.19 0.18 0.19
noisy 0.38 0.36 0.39 0.15 0.13 0.16 8.20 7.93 8.72 0.20 0.18 0.21
shrec 0.52 0.48 0.59 0.28 0.24 0.35 5.53 4.51 6.20 0.17 0.16 0.19
9 faket 0.36 0.31 0.39 0.14 0.11 0.16 8.70 7.89 9.76 0.20 0.19 0.21
noisy 0.37 0.35 0.38 0.15 0.13 0.15 8.38 8.10 8.92 0.19 0.17 0.20
shrec 0.57 0.50 0.61 0.34 0.26 0.38 4.72 4.15 5.78 0.14 0.13 0.15
Table S1: Standard metrics, related to Evaluation. Despite evaluating the quality of our simulations directly on downstream tasks (particle localization and classification), we also compute standard metrics. Mean Absolute Error (MAE), Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) metrics for simulated tilt-series are computed in contrast to their noiseless counterparts. Since SSIM assumes non-negative images with a well-defined range of values, we normalize each tilt-series to interval $[0,1]$. Then we compute the MAE, MSE, PSNR, and SSIM per tilt and provide the mean value of each metric per tilt-series along with its $10^{th}$ and $90^{th}$ percentiles.
Figure S1: Visualization of the inputs (Noiseless, Style), output (FakET) and ground truth (SHREC) projections of the first 3 training tomograms at 0° and 40°. Style is always taken from the $(N+1)^{th}$ training tomogram. To minimize the impact of random extreme values (outliers) on contrast, we visualize the Style, FakET, and SHREC projections within the $1^{st}$ and $99^{th}$ percentiles of the pixel intensity. The percentiles are computed based on the whole tilt-series, not just one tilt. We chose the viridis over greyscale colormap for its perceptual uniformity, accessibility for color vision deficiencies, and enhanced detail visibility. Colorbar indicates simulated intensities in arbitrary units.
Figure S2: Visualization of the inputs (Noiseless, Style), output (FakET) and ground truth (SHREC) projections of the $4^{th}$ - $6^{th}$ training tomograms at 0° and 40°. Style is always taken from the $(N+1)^{th}$ training tomogram. To minimize the impact of random extreme values (outliers) on contrast, we visualize the Style, FakET, and SHREC projections within the $1^{st}$ and $99^{th}$ percentiles of the pixel intensity. The percentiles are computed based on the whole tilt-series, not just one tilt. We chose the viridis over greyscale colormap for its perceptual uniformity, accessibility for color vision deficiencies, and enhanced detail visibility. Colorbar indicates simulated intensities in arbitrary units.
Figure S3: Visualization of the inputs (Noiseless, Style), output (FakET) and ground truth (SHREC) projections of the last 3 training tomograms at 0° and 40°. Style is always taken from the $(N+1)^{th}$ training tomogram. To minimize the impact of random extreme values (outliers) on contrast, we visualize the Style, FakET, and SHREC projections within the $1^{st}$ and $99^{th}$ percentiles of the pixel intensity. The percentiles are computed based on the whole tilt-series, not just one tilt. We chose the viridis over greyscale colormap for its perceptual uniformity, accessibility for color vision deficiencies, and enhanced detail visibility. Colorbar indicates simulated intensities in arbitrary units.